# R-SFLLM: Jamming Resilient Framework for Split Federated Learning with Large Language Models

### v2

@@ II-B1 Assumptions on the Loss Function
$L$ is $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smooth coordinate-wisely, i.e. there exist coefficient vectors $\boldsymbol{L_{0}},\boldsymbol{L_{1}}\in\mathbb{R}^{E}$ such that for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{E}$ with $\|\boldsymbol{x}-\boldsymbol{y}\|\leq\frac{1}{\|\boldsymbol{L_{1}}\|_{\infty}}$ it holds for all $j\in[E]=\{1,\ldots,E\}$ that
$\left|\frac{\partial L(\boldsymbol{y})}{\partial x_{j}}-\frac{\partial L(\boldsymbol{x})}{\partial x_{j}}\right|\leq\left(\frac{L_{0,j}}{\sqrt{E}}+L_{1,j}\left|\frac{\partial L(\boldsymbol{x})}{\partial x_{j}}\right|\right)\cdot||\boldsymbol{y}-\boldsymbol{x}||_{2}.$
@@
$||\nabla^{2}L(\boldsymbol{x})||\leq L_{0}+L_{1}||\nabla L(\boldsymbol{x})||.$
The coordinate-wise $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smoothness implies that smoothness may vary for each coordinate of the input space. This particularly pertains to LLMs as it has been shown in [32] that variance can be observed across mostly every transformer layer, such that each layer coordinate $j$ satisfies an own $(L_{0,j},L_{1,j})$ pair. Thus, if the coefficients $L_{1,j}$ are non-zero, smoothness is potentially unbounded. In contrast, if all $L_{1,j}$ are strictly zero, the original Lipschitz smoothness is recovered. In [32], the following Lemma has been established, relating the coordinate-wise smoothness to the loss divergence.
@@
[32] Let $L$ be $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smooth coordinate-wisely. Then for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{E}$ with $||\boldsymbol{x}-\boldsymbol{y}||_{2}\leq\frac{1}{||\boldsymbol{L_{1}}||_{\infty}}$, we have
@@
$+\sum_{j=1}^{E}\frac{\left(\frac{L_{0,j}}{\sqrt{E}}+L_{1,j}\left|\frac{\partial L(\boldsymbol{x})}{\partial x_{j}}\right|\right)||\boldsymbol{y}-\boldsymbol{x}||_{2}}{2}\ |y_{j}-x_{j}|.$ (11)
@@ II-B2 Upper Bound on the LLM Loss Divergence
In (12), the loss gradient $\|\nabla L(\boldsymbol{x})\|_{2}$ might be unbounded, particularly when several coordinates need to be considered. However, common practice in deep learning suggests to bound gradients manually by means of gradient clipping [33] using a clipping threshold $\tau>0$, thereby preventing exploding gradients, i.e.
@@ II-B3 Relating the Model Error to the Communication MSE
Let $\boldsymbol{e}_{q},\hat{\boldsymbol{e}}_{q}\in\mathbb{R}^{E}$ be the true and corrupted word embeddings and let $s_{qnk},\hat{s}_{qnk}\in\mathbb{C}$ be the corresponding true and corrupted transmit symbols. Then, it holds that
@@
Let $L:\mathbb{R}^{E}\rightarrow\mathbb{R}$ satisfy Assumptions 1 and 2 with coordinate-wise smoothness parameters $\boldsymbol{L}_{0},\boldsymbol{L}_{1}\in\mathbb{R}^{E}$ and let $\boldsymbol{e}_{q},\hat{\boldsymbol{e}}_{q}\in\mathbb{R}^{E}$ be the true and corrupted word embeddings with $\|\boldsymbol{e}_{q}-\hat{\boldsymbol{e}}_{q}\|\leq\frac{1}{\|\boldsymbol{L_{1}}\|_{\infty}}$. Then, it holds that
@@ II-B4 Practical Interpretation of Results
Proposition 2 provides an upper bound on the model error, defined by the expected loss divergence between legitimate and corrupted embeddings, which is directly dependent on the MSE of the wireless communication system. Therein, the proximity condition $\|\boldsymbol{e}_{q}-\hat{\boldsymbol{e}}_{q}\|_{2}\leq\frac{1}{\|\boldsymbol{L_{1}}\|_{\infty}}$ sets a practical constraint on the distance between legitimate and jammed embeddings, which needs to be small enough for the smoothness condition to hold. As outlined in [32], this ensures the stability of the gradient behavior, preventing numerical instabilities and unreliable approximations as gradient-dependent optimization algorithms might struggle to converge. By applying gradient clipping, we ensure that $\|\nabla_{\boldsymbol{e}_{q}}L(\boldsymbol{e}_{q})\|_{2}$ is bounded by $\tau$, thereby stabilizing the training process as the assumptions underlying the optimization methods are not violated. In particular, Corollary 2 ensures that the model error does not explode and is only dependent on the $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smoothness coefficients, the clipping threshold $\tau$ and the communication MSE $\mathbb{E}[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}]$, independent of whether the proximity condition is fulfilled or not. In the context of jamming, this allows for the consideration of arbitrary adversaries, including worst-case scenarios. Hence, even if the resulting jammed word embeddings do not fulfill the proximity condition, for example due to excessive noise or sophisticated attack strategies that may flip the embedding label, our analysis remains applicable, aligning with best practices in deep learning. This new insight thus establishes a formal relationship between the transformer-based LLM architecture, its semantic word embeddings, and the wireless medium, thereby emphasizing the importance of the wireless communication system and its resilience to adversarial jamming in the quality of distributed training. To the best of our knowledge, this is the first formal characterization of such a relationship for practical DCML applications, such as SFL, with LLMs over wireless networks. In particular, we consider a generalized smoothness assumption on the loss function, which is often omitted in previous works but required for a proper analysis. This assumption explains why techniques such as gradient clipping work and may be necessary during training, and how corrupted model inputs affect the loss divergence, a critical measure for the ML training performance. Consequently, as the model error directly depends on the communication MSE, a wireless approach to resilience in SFL is not only justified but required, thus instructing us to maximize the SINR.
II-C Minimum System Rate for Reliable SFL with LLMs
@@
To characterize the minimum system rate, under which the communication link can support SFL reliably, we need to identify outage conditions caused by jamming. To this end, we provide three remarks. In Remark 1, we first derive a lower bound for the per resource allocation symbol error. We use this lower bound in Remark 2 to establish a general lower bound on the outage rate which is dependent on the $\boldsymbol{L}_{1}$ constant from Lemma 1. In Remark 3, we derive the outage rate for the particular case where Proposition 2 is fulfilled with equality.
@@
$\mathbb{E}\left[\left|L(\boldsymbol{e}_{q})-L(\hat{\boldsymbol{e}}_{q})\right|\right]=\epsilon_{q}\leq\gamma_{q}\sqrt{\mu_{q}}+\delta_{q}\mu_{q}$ (29)
where $\epsilon_{q},\gamma_{q},\delta_{q}$ are substitutes for the corresponding expressions in (20) and $\mu_{q}$ as in (25) from Remark 1. Now, we require $\epsilon_{q}=\gamma_{q}\sqrt{\mu_{q}}+\delta_{q}\mu_{q}$ with equality. By setting $y=\sqrt{\mu_{q}}=\sqrt{r_{q}\exp\{\-R_{q}/r_{q}\}}$, we can equivalently state that
@@
Thus, we obtain $R_{\textrm{out},2}$ by solving $y^{2}=r_{q}\cdot\exp\{R_{q}/r_{q}\}$ for $R_{q}$ as
@@
From Remarks 2 and 3, we can conclude that both the communication system and SFL perform reliably for system rates $R>R_{\textrm{out}}=\min\{R_{\textrm{out},1},R_{\textrm{out},2}\}$. Note that the minimum rate $R_{\textrm{out}}$ depends on either $\boldsymbol{L}_{1}$ from to the proximity condition or on the particular realization of the adversarial jammer captured by the MSE. Thus, particularly strong or worst-case jammers may be able to considerably decrease the system rate below $R_{\textrm{out}}$, such that neither automatic repeat requests (ARQs) nor other upper-layer resilience mechanisms can be applied. This advocates the need for proactive, resilient-by-design physical layer solutions to anti-jamming to ensure robust and reliable wireless SFL, even under worst-case conditions where reactive network defenses cannot suffice. Moreover, the proximity condition involving $\boldsymbol{L}_{1}$ further indicates that the system can still maintain robustness if the jammer introduces only moderate noise, thereby not violating smoothness as long as the embeddings are not too divergent. This insight further opens up the door for adversarial training aspects under which robustness can even be increased due to controlled noise exposure [24], which we verify in Section IV.
@@ III R-SFLLM Anti-Jamming Framework
III-A Anti-Jamming Strategy and Optimization Problem
As a result of Proposition 2, jammed SFL word embeddings $\hat{\boldsymbol{e}}_{q}$ lead to a deviation from the ground truth in the deterministic loss function $L$. Thus, the anti-jamming objective can be generally formulated as the minimization of the expected loss divergence, i.e. $\min\ \mathbb{E}\left[\left|L(\boldsymbol{e}_{q})-L(\hat{\boldsymbol{e}}_{q})\right|\right]$. Using Corollary 2, we can instead minimize $J(\boldsymbol{s}_{q},\hat{\boldsymbol{s}}_{q})=\sqrt{\mathbb{E}\left[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}\right]}+\mathbb{E}\left[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}\right]$, which is only dependent on the MSE and where we imply using gradient clipping during training. This problem can be equivalently interpreted as the maximization of the SINR for each user $q\in\mathcal{Q}$, respectively, or more generally, as the maximization of the sum rate. To this end, we derive an expression for the achievable sum rate as follows:
@@ III-A Anti-Jamming Strategy and Optimization Problem
In order to incorporate anti-jamming in SFL by-design, we need to jointly optimize over beamforming, user scheduling, and resource allocation constraints, thus introducing resilience proactively at the bit level. To this end, we pose the following optimization problem that is applied for all $q\in\mathcal{Q},n\in\mathcal{N}$, and $k\in\mathcal{K}$:
[…]
@@
In this problem, constraints (36a) to (36c) incorporate the binary user scheduling, power allocation and resource block limitation, with $B_{q}=|\mathcal{R}_{q}|$ being the maximum number of resource allocation blocks for each user $q$. Constraint (36d) captures the power budget and constraint (36e) ensures the unit normalization of the beamforming vector. However, due to the binary user scheduling variable $\alpha_{qnk}$, the optimization is a non-linear, non-convex mixed-integer problem and thus NP-hard in general. In addition to NP-hardness, the sum rate further depends on the composite noise covariance matrix $\boldsymbol{C}_{\boldsymbol{z}_{nk}}$. Since the latter characterizes the adversarial jamming strategy, which is not known to the legitimate SFL parties, a surrogate expression $\widetilde{\boldsymbol{C}}_{\boldsymbol{z}_{nk}}$ needs to be found, which effectively approximates the true covariance, i.e. $\widetilde{\boldsymbol{C}}_{\boldsymbol{z}_{nk}}\approx\boldsymbol{C}_{\boldsymbol{z}_{nk}}$.
III-B Role of Sensing-Assisted Jamming DoA Information
@@
By inserting (37) into (34), we hence obtain a lower bound on $R$. Note that in general $\eta$ is unknown since it depends on the unknown jamming setup. Thus, we consider it as a hyperparameter, which controls the resilience level of our system. In [34], we showed that by choosing $\eta$ to be much larger than the noise level $\sigma^{2}$, i.e. $\eta\gg\sigma^{2}$, we coincide with the case where the jammer setup is known, that is where $\gamma\approx\tilde{\gamma}$. In this case, the SINR can be maximized by maximizing the lower bound $\tilde{\gamma}$ instead. Thus, $\widetilde{\boldsymbol{C}}_{\boldsymbol{z}_{nk}}$ constitutes a conservative approximation of $\boldsymbol{C}_{\boldsymbol{z}_{nk}}$, ensuring it does not underestimate the impact of noise and adversarial jamming, indicated by the Löwner order $\succeq$ in (37). We carefully validated this conjecture in [22] for $\eta=10$ in several jamming scenarios for a range of power budgets $P_{J}<\infty$. Consequently, the availability of the DoAs alleviates the need to know the exact jamming statistics.
@@
4Initialize the interference-plus-noise covariance matrix to $\boldsymbol{X}_{qnk}=\boldsymbol{C}_{\boldsymbol{z}_{nk}}$

### v3

@@ II-B1 Assumptions on the Loss Function
$L$ is $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smooth coordinate-wisely, i.e. there exist coefficient vectors $\boldsymbol{L_{0}},\boldsymbol{L_{1}}\in\mathbb{R}^{E}$ such that for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{E}$ with $\|\boldsymbol{x}-\boldsymbol{y}\|\leq\frac{1}{\|\boldsymbol{L_{1}}\|_{\infty}}$ it holds $\forall j\in[E]$ that
[…]
@@
[…]
The coordinate-wise $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smoothness implies that smoothness may vary for each coordinate of the input space. This particularly pertains to LLMs as it has been shown in [38] that variance can be observed across mostly every transformer layer, such that each layer coordinate $j$ satisfies an own $(L_{0,j},L_{1,j})$ pair. Thus, if the coefficients $L_{1,j}$ are non-zero, smoothness is potentially unbounded. In contrast, if all $L_{1,j}$ are strictly zero, the original Lipschitz smoothness is recovered. In [38], the following Lemma has been established, relating the coordinate-wise smoothness to the loss divergence:
@@
[…]
@@
[…]
@@ II-B2 Upper Bound on the LLM Loss Divergence
In (12), the gradient loss $\|\nabla L(\boldsymbol{x})\|_{2}$ might be unbounded, particularly when several coordinates need to be considered. However, common practice in deep learning suggests to bound gradients manually via gradient clipping [39] using a clipping threshold $\tau>0$, thereby preventing exploding gradients, i.e.
@@ II-B3 Relating the Model Error to the Communication MSE
Let $\boldsymbol{e}_{q},\hat{\boldsymbol{e}}_{q}\in\mathbb{R}^{E}$ be the true and corrupted embeddings and let $s_{qnk},\hat{s}_{qnk}\in\mathbb{C}$ be the corresponding true and corrupted transmit symbols. Then, it holds that
@@
Let $L:\mathbb{R}^{E}\rightarrow\mathbb{R}$ satisfy Assumptions 1 and 2 with coordinate-wise smoothness parameters $\boldsymbol{L}_{0},\boldsymbol{L}_{1}\in\mathbb{R}^{E}$, and let $\boldsymbol{e}_{q},\hat{\boldsymbol{e}}_{q}\in\mathbb{R}^{E}$ be the true and corrupted embeddings with $\|\boldsymbol{e}_{q}-\hat{\boldsymbol{e}}_{q}\|\leq\frac{1}{\|\boldsymbol{L_{1}}\|_{\infty}}$. Then, it holds that
@@ II-B4 Practical Interpretation of Results
Proposition 2 provides an upper bound on the model error, defined by the expected loss divergence between legitimate and corrupted embeddings, which is directly dependent on the MSE of the wireless communication system. Therein, the proximity condition $\|\boldsymbol{e}_{q}-\hat{\boldsymbol{e}}_{q}\|_{2}\leq\frac{1}{\|\boldsymbol{L_{1}}\|_{\infty}}$ sets a practical constraint on the distance between legitimate and jammed embeddings, which needs to be small enough for the smoothness condition to hold. As outlined in [38], this ensures the stability of the gradient behavior, preventing numerical instabilities and unreliable approximations as gradient-dependent optimization algorithms might struggle to converge. By applying gradient clipping, we ensure that $\|\nabla_{\boldsymbol{e}_{q}}L(\boldsymbol{e}_{q})\|_{2}$ is bounded by $\tau$, thereby stabilizing the training process as the assumptions underlying the optimization methods are not violated. In particular, Corollary 2 ensures that the model error does not explode and is only dependent on the $(\boldsymbol{L_{0}},\boldsymbol{L_{1}})$-smoothness coefficients, the clipping threshold $\tau$, and the communication MSE $\mathbb{E}[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}]$, independent of whether the proximity condition is fulfilled or not. In the context of jamming, this allows for the consideration of arbitrary adversaries, including worst-case scenarios. Hence, even if the resulting jammed embeddings do not fulfill the proximity condition, for example due to excessive noise or sophisticated attack strategies that might flip the embedding label, our analysis remains applicable, aligning with best practices in deep learning. This new insight establishes a formal relationship between the transformer-based LLM architecture, its semantic embeddings, and the wireless medium, thereby emphasizing the importance of the wireless communication system and its resilience to adversarial jamming in the quality of distributed training. To the best of our knowledge, this is the first formal characterization of such a relationship for practical DCML with LLMs over wireless networks. In particular, we consider a generalized smoothness assumption on the loss function, which is often omitted in previous works but required for a proper analysis. This assumption explains why techniques such as gradient clipping work and might be necessary during training, and how corrupted model inputs affect the loss divergence, a critical measure for the ML training performance. Consequently, as the model error directly depends on the communication MSE, a wireless approach to resilience in SFL is not only justified but required, thus instructing us to maximize the SINR.
II-C Minimum System Rate for Reliable SFL with LLMs.
@@
To characterize the minimum system rate under which the communication link can support SFL reliably, we need to identify outage conditions caused by jamming. To this end, we provide three remarks. In Remark 1, we first derive a lower bound for the per resource allocation symbol error. We use this lower bound in Remark 2 to establish a general lower bound on the outage rate, which is dependent on the $\boldsymbol{L}_{1}$ constant from Lemma 1. In Remark 3, we derive the outage rate for the particular case where Proposition 2 is fulfilled with equality.
@@
$\mathbb{E}\left[\left|L(\boldsymbol{e}_{q})-L(\hat{\boldsymbol{e}}_{q})\right|\right]=\epsilon_{q}\leq\gamma_{q}\sqrt{\mu_{q}}+\delta_{q}\mu_{q},$ (29)
where $\epsilon_{q},\gamma_{q},\delta_{q}$ are substitutes for the corresponding expressions in (20), and $\mu_{q}$ as in (25) from Remark 1. Now, we require $\epsilon_{q}=\gamma_{q}\sqrt{\mu_{q}}+\delta_{q}\mu_{q}$ with equality. By setting $y=\sqrt{\mu_{q}}=\sqrt{r_{q}\exp\{\-R_{q}/r_{q}\}}$, we can equivalently state that
@@
Thus, we obtain $R_{\textrm{out},2}$ by solving $y^{2}=r_{q}\cdot\exp\{\frac{R_{q}}{r_{q}}\}$ for $R_{q}$:
@@
From Remarks 2 and 3, we can conclude that both the communication system and SFL perform reliably for system rates $R>R_{\textrm{out}}=\min\{R_{\textrm{out},1},R_{\textrm{out},2}\}$. Note that the minimum rate $R_{\textrm{out}}$ depends on either $\boldsymbol{L}_{1}$ from to the proximity condition or on the particular realization of the adversarial jammer captured by the MSE. Thus, particularly strong or worst-case jammers may be able to considerably decrease the system rate below $R_{\textrm{out}}$, such that neither automatic repeat requests (ARQs) nor other upper-layer resilience mechanisms can be applied. As $R_{\textrm{out}}$ scales with the number of users and available bandwidth, robust anti-jamming and resource allocation become essential for sustaining performance in larger SFL deployments. To preserve SFL training under diverse and unpredictable disruptions, we specifically need to demand more than mere robustness, which only accounts for moderate or known distribution shifts, and instead require resilient defenses. This advocates the need for proactive, resilient-by-design physical layer solutions to anti-jamming to ensure resilient and reliable wireless SFL, even under worst-case conditions where reactive network defenses cannot suffice. Moreover, the proximity condition involving $\boldsymbol{L}_{1}$ further indicates that the system can still maintain resilience if the jammer introduces only moderate noise, thereby not violating smoothness as long as the embeddings are not too divergent. This insight further opens up the door for adversarial training aspects under which resilience can even be increased due to controlled noise exposure [28], which we verify in Section IV.
@@ III R-SFLLM Anti-Jamming Framework
III-A Anti-Jamming Strategy and Optimization Problem.
As a result of Proposition 2, jammed embeddings $\hat{\boldsymbol{e}}_{q}$ lead to a deviation from the ground truth in the deterministic loss function $L$. Thus, the anti-jamming objective can be generally formulated as the minimization of the expected loss divergence, i.e. $\min\ \mathbb{E}\left[\left|L(\boldsymbol{e}_{q})-L(\hat{\boldsymbol{e}}_{q})\right|\right]$. Using Corollary 2, we can instead minimize $J(\boldsymbol{s}_{q},\hat{\boldsymbol{s}}_{q})=\sqrt{\mathbb{E}\left[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}\right]}+\mathbb{E}\left[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}\right]$, which is only dependent on the MSE and where we imply using gradient clipping during training. This problem can be equivalently interpreted as the maximization of the SINR for each user $q\in\mathcal{Q}$, respectively, or more generally, as the maximization of the sum rate. To this end, we derive an expression for the achievable sum rate as follows:
@@
To incorporate anti-jamming in SFL by-design, we need to jointly optimize over beamforming, user scheduling, and resource allocation constraints, thus introducing resilience proactively at the bit level. To this end, we pose the following optimization problem applied for $q\in\mathcal{Q},n\in\mathcal{N}$, and $k\in\mathcal{K}$:
$\max_{\begin{subarray}{c}\alpha_{qnk},p_{qnk},\boldsymbol{w}_{qnk},\boldsymbol{v}_{qnk}\\ (n,k)\in\mathcal{R}_{q},q\in\mathcal{Q}\end{subarray}}\quad$ $R\quad\text{s.t.}\quad$ (36)
@@
In this problem, constraints (36a) to (36c) incorporate the binary user scheduling, power allocation and resource block limitation, with $B_{q}=|\mathcal{R}_{q}|$ being the maximum number of resource allocation blocks for each user $q$. Constraint (36d) captures the power budget and constraint (36e) ensures the unit normalization of the beamforming vector. However, due to the binary user scheduling variable $\alpha_{qnk}$, the optimization is a non-linear, non-convex mixed-integer problem and thus NP-hard in general. In addition to NP-hardness, the sum rate further depends on the composite noise covariance matrix $\boldsymbol{C}_{\boldsymbol{z}_{nk}}$. Since the latter characterizes the adversarial jamming strategy, which is not known to the legitimate SFL parties, a surrogate expression $\widetilde{\boldsymbol{C}}_{\boldsymbol{z}_{nk}}$ needs to be found that effectively approximates the true covariance, i.e. $\widetilde{\boldsymbol{C}}_{\boldsymbol{z}_{nk}}\approx\boldsymbol{C}_{\boldsymbol{z}_{nk}}$.
III-B Role of Sensing-Assisted Jamming DoA Information.
@@
By inserting (37) into (34), we hence obtain a lower bound on $R$. Note that in general $\eta$ is unknown since it depends on the unknown jamming setup. Thus, we consider it as a conscious hyperparameter, which controls the resilience level of our system. In [40], we showed that by choosing $\eta$ to be much larger than the noise level $\sigma^{2}$, i.e. $\eta\gg\sigma^{2}$, we coincide with the case where the jammer setup is known, that is where $\gamma\approx\tilde{\gamma}$. In this case, the SINR can be maximized by maximizing the lower bound $\tilde{\gamma}$ instead. Thus, $\widetilde{\boldsymbol{C}}_{\boldsymbol{z}_{nk}}$ constitutes a conservative approximation of $\boldsymbol{C}_{\boldsymbol{z}_{nk}}$, ensuring it does not underestimate the impact of noise and adversarial jamming, indicated by the Löwner order $\succeq$ in (37). We carefully validated this conjecture in [22] for $\eta=10$ in several jamming scenarios for a range of power budgets $P_{J}<\infty$. Consequently, the availability of the DoAs alleviates the need to know the exact jamming statistics. In future 6G networks, having access to jamming DoAs is a realistic assumption, as native wireless sensing services can provide this information, for example, as part of ISAC and RIS protocols. We refer the interested reader to a comprehensive overview of such protocols in [33] and [34].
@@
4
5Initialize the interference-plus-noise covariance matrix to $\boldsymbol{X}_{qnk}=\boldsymbol{C}_{\boldsymbol{z}_{nk}}$